### v1

@@ Abstract
Objective: This study proposes a Natural Language Processing pipeline to evaluate prescription appropriateness in Italian referrals, where reasons for prescriptions are recorded only as free text, complicating automated comparisons with guidelines. The pipeline aims to derive, for the first time, a comprehensive summary of the reasons behind these referrals and a quantification of their appropriateness. While demonstrated in a specific case study, the approach is designed to generalize to other types of examinations.
Methods: Leveraging embeddings from a transformer-based model, the proposed approach clusters referral texts, maps clusters to labels, and aligns these labels with existing guidelines. We present a case study on a dataset of 496,971 referrals, consisting of all referrals for venous echocolordopplers of the lower limbs between 2019 and 2021 in the Lombardy Region. A sample of 1,000 referrals was manually annotated to validate the results.
Results: The pipeline exhibited high performance for referrals’ reasons (Prec=92.43%, Rec=83.28%) and excellent results for referrals’ appropriateness (Prec=93.58%, Rec=91.52%) on the annotated subset. Analysis of the entire dataset identified clusters matching guideline-defined reasons - both appropriate and inappropriate - as well as clusters not addressed in the guidelines. Overall, 34.32% of referrals were marked as appropriate, 34.07% inappropriate, 14.37% likely inappropriate, and 17.24% could not be mapped to guidelines.
Conclusions: The proposed pipeline effectively assessed prescription appropriateness across a large dataset, serving as a valuable tool for health authorities. Findings have informed the Lombardy Region’s efforts to strengthen recommendations and reduce the burden of inappropriate referrals.
@@
Natural language processing, Referrals, Prescription appropriateness, Public healthcare system, Quality of healthcare
1 Background and significance
@@
In Italy, the National Health System (NHS) is the primary provider of healthcare services, offering universal coverage, with decentralized administration at the regional level [1]. Despite a strong performance in international rankings, the Italian NHS has faced growing challenges in providing timely access to healthcare services in recent years, leading many citizens to turn to private providers [2].
@@
One significant contributor to these delays is the issue of prescription appropriateness, a concern shared by many advanced healthcare systems [3, 4]. Inappropriate prescriptions have a threefold negative impact on the healthcare system: waste of resources, increased waiting times, and reduced effectiveness of the care provided to citizens, who face delays in obtaining accurate diagnoses [5]. In the Italian NHS, referrals from general practitioners (GPs) or specialized physicians are necessary to book examinations. Assessing prescription appropriateness requires understanding the reasons underlying these referrals.
When Italian physicians fill out a referral, they have to use a standardized form, providing the reason for the prescription in a specific free-text field called Clinical Question. The unstructured nature of this field, coupled with the lack of standardized alternatives, makes it challenging to analyze. Addressing this issue requires Natural Language Processing (NLP) techniques.
@@
The number of studies analyzing texts of Italian referrals is very limited. Venturelli et al. [6] used proprietary commercial software (Clinika VAP [7]), for which the NLP models or algorithms are not disclosed. This software provides an appropriate/inappropriate label for each referral, enabling the quantification of inappropriate referrals. However, it does not provide a clustered analysis of the referrals, with a quantification of the different prescription reasons. To the best of our knowledge, the only other study applying NLP on Italian referrals’ texts is [8], which focused on extracting and analyzing follow-up waiting times from clinical questions to investigate timely access to care.
While several studies have evaluated prescription appropriateness for medications within the Italian NHS using structured data or manual text analysis (e.g., [9, 10, 11]), research on the appropriateness of examinations is more limited [12, 13] and has not incorporated natural language processing.
@@
A few studies developed decision-support systems for guideline adherence, both in primary care [14, 15] and hospital settings [16, 17]. These tools assist physicians through closed-form questions, possibly complemented by keyword matching on the text they write. While effective in specific scenarios [18, 19], such tools are not suited for retrospective analyses of their performance or for quantifying the prevalence of appropriate or inappropriate referrals.
@@
Beyond the Italian context, Haddad et al. [20] analyzed Russian referrals’ texts for prescription appropriateness using labelled data; Sagheb et al. [21] designed a rule-based algorithm to verify guidelines adherence for asthma patients from EHR documents; similarly in [22] a rule-based approach was applied to assess guideline adherence for gout patients. Dimitrov et al. [23] conducted a nationwide evaluation of breast cancer guideline adherence in EHRs using proprietary commercial software (Danny platform [24]), which incorporated rules and manual reviews. Villeana et al. [25], analyzed Chilean referrals’ texts using labelled data to binary classify them for correct priority classes.
These studies, however, are often disease-specific or reliant on labelled data or proprietary software, limiting their broader applicability.
2 Objective
This study proposes an NLP-based pipeline to assess the appropriateness of Italian referrals, developing a model designed for unsupervised analysis across different types of medical examinations.
In 2015, the Lombardy Region, the most populated Italian region with approximately 10 million inhabitants - akin to a small state - issued two resolutions (DGRs) focusing on prescription appropriateness in angiology [26] and gastroenterology [27], two fields where the literature reports high rates of inappropriate prescriptions, both in Italy and internationally [28, 29, 30, 31, 32]. These resolutions established guidelines specifying appropriate and inappropriate clinical questions for a variety of examinations.
Building on these efforts, we analyzed examinations covered by these resolutions, contributing to the issuance of a new resolution in December 2023 that reinforced the previous guidelines [33]. This paper focuses on a case study of one of these examinations: the venous echocolordoppler of the lower limbs. In particular, we analyzed a dataset consisting of all referrals prescribed in the Lombardy Region for this exam between 2019 and 2021, for a total of 496,971 referrals, encompassing 229,509 unique clinical questions.
The analysis presented in this study aims to verify the effectiveness of the 2015 resolutions in enforcing prescriptive appropriateness, providing for the first time an overview of the main reasons why these examinations are prescribed, with an approximate quantification of the number of referrals for each disease/symptom and their alignment on existing guidelines, where possible.
Given the absence of labelled data and the need to uncover reasons not addressed in guidelines, we framed this as a clustering problem. In particular, our pipeline clusters clinical questions by exploiting the embeddings produced by a fine-tuned BERT model [34]. These clusters are summarized and, where possible, mapped to guidelines to determine whether they are appropriate, inappropriate, or not covered by the guidelines.
A subset of 1,000 distinct clinical questions was manually annotated to validate the pipeline. Results are reported for this validation subset to verify the correctness of the pipeline output and for the full dataset to analyze prescription appropriateness at the regional level, including stratifications by year (pre-, during, and post-COVID pandemic), geographical area, and type of prescribing physician. Additional results for another type of examination in the gastroenterological domain are included in Appendix D.
The main contributions of this study are the following:
  1. 1.
    The first open-source, fully unsupervised pipeline for assessing prescription appropriateness from Italian referrals, applicable to any type of examination.
  2. 2.
    The first region-wide analysis in Italy of referrals’ reasons and their appropriateness.
3 Material and Methods
@@ 3.1 Data
3.1.1 Referrals dataset
The dataset analyzed in this study is extracted from all referrals for venous echocolordoppler of the lower limbs prescribed in the Lombardy Region between 2019 and 2021 (496,971 referrals). Each referral includes the clinical question, the year, the Local Health Authority (LHA) of the prescribing physician and the type of prescribing physician. Physician types are categorized as general practitioners (GPs) or specialized physicians, while LHAs represent the eight healthcare administrative areas within the Lombardy Region. Figure 1 illustrates the distribution of referrals among these groups.
The analysis focuses on the clinical question, a free-text field typically containing one or few sentences. These texts exhibit frequent spelling and typing errors, heavy use of abbreviations, inconsistent punctuation, and non-standard capitalization. The dataset includes 229,509 distinct clinical questions, but only 15% appear more than once. The median length is 71 characters (IQR=39). Only 189 (0.08%) contain a meaningful ICD code, underscoring the necessity of employing NLP for this analysis.
The clinical questions should not contain personal information about patients or physicians. A preliminary check was performed by the Lombardy Region IT provider, masking any suspect personal information before data extraction.
@@ 3.1.2 Manual annotations
To validate the identified clusters, a randomly selected subset of 1,000 distinct clinical questions was manually labelled. The labels correspond to categories representing diseases or symptoms that prompted the referral. During an initial review of this subset, ambiguous cases were identified, and an angiologist was consulted to solve doubts and align clusters with prescription guidelines. Annotation guidelines were subsequently refined, and the entire subset was annotated accordingly. Two other reviews of the annotations corrected remaining errors and clarified uncertainties, ensuring the final labelled dataset was error-free. More details on the annotations are provided in Appendix A.
@@ 3.1.3 Prescription guidelines
The guidelines used as a reference to assess referral appropriateness are extracted from [26]. In particular, we extracted the appropriate and inappropriate diseases and symptoms for the venous echocolordoppler of the lower limbs. An English translation is reported in Figure 2.
Figure 1: Referrals distributions across years (A), type of prescribing physicians (B) and LHA (C)
Figure 2: Guidelines for referrals for venous echocolordoppler of the lower limbs in the Lombardy Region. Translated from the Appendix 3 of [26].
3.2 Pipeline for prescription appropriateness
@@
Figure 3 depicts blocks composing our pipeline, which are detailed below.
@@ 3.2.1 Pre-processing
Physicians write referrals during patients’ examinations, with limited time available. This leads them to write clinical questions that are far from well-readable and correct texts. The grammar of the sentences is often poor, there are frequent typing errors, and the punctuation is either not used or used incorrectly. The use of uppercase letters does not have any particular meaning, and the use of abbreviations is very frequent. Even in the single type of examination analyzed in this study, multiple abbreviations can indicate the same disease/symptoms. In the manually labelled subset, 15% of clinical questions were marked as having typing errors and 68% as containing abbreviations.
To address these issues, suitable pre-processing steps were applied:
@@
We removed all punctuation and lowercase sentences due to the inconsistent use of these elements.
[…]
Examples of pre-processing effects are reported in Appendix C.
@@ 3.2.2 Text embedding
To appropriately cluster the clinical questions, it is necessary to produce a representation capturing the central semantics of these texts without getting confused by their widely present noise. Because of this, vocabulary-based representations, such as TF-IDF [35], cannot be expected to work well for this task. The main alternatives are static and contextual neural network-based embeddings.
@@
On the contrary, context-based embeddings produced by transformer models consider the surrounding context while embedding each token, and have often shown superior performances in the recent clinical NLP literature [37, 38, 39, 40]. The problem is particularly challenging in our case due to the limited number of embedding models available for the Italian language, especially for the clinical domain.
To this purpose, the same model used in [41] was adopted to cluster sentences extracted from Italian referrals, and we verified that even with clinical questions, its performances are superior to other models. This model consists of a fine-tuned version of the last Italian version of BERT, Umberto [42], on a dataset of publicly available medical documents in Italian and other languages (translated into Italian) [43].
@@ 3.2.3 Clustering
HDBscan was adopted to cluster the embeddings since it does not require eliciting in advance the number of clusters and it allows a not complete assignment of items to clusters. HDBscan requires defining the minimum size of the clusters to identify. We verified that the best results are achieved when this parameter is progressively reduced while iterating the clustering on the previously unclustered data. This approach reduced the number of unclustered points, addressing the varying cluster sizes in our dataset.
An additional issue in the clustering procedure was related to the size of our dataset. On a machine with 32 GB of RAM, it was not possible to run the clustering algorithm on the whole dataset. Because of this, we had to divide it into chunks of 30,000 clinical questions and cluster them separately. The subsequent summarization and mapping steps merged the clusters found in different chunks and corresponding to the same diseases/symptoms.
@@ 3.2.4 Cluster summarization
Summarizing the identified clusters is necessary to understand the symptoms or diseases corresponding to each cluster. To automatize this step, we extracted keywords from each cluster. The most frequent words in each cluster were initially selected, considering a decreasing frequency threshold until at least two keywords were found. Among these, we excluded those with high frequency in at least half of the clusters. This allowed us to exclude not only stop words but also frequent words for the specific exam (e.g., leg, lower limbs, right, left in our case study).

### v2

@@ Abstract
Background: Assessing the appropriateness of diagnostic referrals is critical for improving healthcare efficiency and reducing unnecessary procedures. However, this task becomes challenging when referral reasons are recorded only as free text rather than structured codes. In the Italian National Health Service (NHS), referral motivations are systematically entered as unstructured text, making large-scale evaluation against official guidelines difficult to automate and scale. To address this gap, we propose a fully unsupervised Natural Language Processing (NLP) pipeline capable of extracting and evaluating referral reasons without relying on labelled datasets, which are costly and time-consuming to produce.
Methods: Our pipeline leverages Transformer-based embeddings pre-trained on Italian medical texts to semantically cluster referral reasons and assess their alignment with appropriateness guidelines. It operates in an unsupervised setting and is designed to generalize across different examination types. We analyzed two complete regional datasets from the Lombardy Region (Italy), covering all referrals between 2019 and 2021 for venous echocolordoppler of the lower limbs (ECD; n=496,971) and flexible endoscope colonoscopy (FEC; n=407,949). The pipeline was developed on the ECD dataset and tested on the FEC dataset. For both, a random sample of 1,000 referrals was manually annotated to measure performance. Results were also stratified by year, region, and physician type.
Results: The pipeline achieved high performance in identifying referral reasons (Precision=92.43% (ECD), 93.59% (FEC); Recall=83.28% (ECD), 92.70% (FEC)) and appropriateness (Precision=93.58% (ECD), 94.66% (FEC); Recall=91.52% (ECD), 93.96% (FEC)). Inappropriate referrals were significantly more prevalent for ECD (34.1%) than for FEC (5.3%). The analysis revealed geographical, temporal, and physician type variations, including a drop in inappropriate referrals during the COVID-19 pandemic in 2020. Additionally, several referral motivations not covered by current guidelines were identified, highlighting areas for policy improvement. These findings contributed to updates to the regional guidelines in Lombardy.
Conclusions: This study presents a robust, scalable, unsupervised NLP pipeline for assessing referral appropriateness from unstructured texts in large, real-world datasets. It demonstrates how such data can be effectively leveraged, providing public health authorities with a deployable AI tool to monitor practices and support evidence-based policy.
@@
Natural language processing, Artificial Intelligence, Machine Learning, Referrals, Unsupervised learning, Examination appropriateness, Guideline adherence, Public healthcare system, Quality of healthcare
pacs:
MSC Classification]68T50
1 Background
@@
In Italy, the National Health Service (NHS) is the primary provider of healthcare services, offering universal coverage, with administration decentralized at the regional level [1]. Despite its strong performance in international rankings, the Italian NHS has faced growing challenges in providing timely access to healthcare services in recent years, with increasing numbers of citizens turning to private providers [2].
@@
One significant contributor to these delays is the issue of referrals appropriateness, a concern shared by many advanced healthcare systems [3, 4]. Inappropriate referrals lead to resource waste, increased waiting times, and delays in accurate diagnoses, ultimately reducing the effectiveness of the healthcare system [5]. Within the Italian NHS, referrals from general practitioners (GPs) and specialists are required to book diagnostic examinations, making it crucial to assess the reasons behind these referrals to ensure their appropriateness.
When issuing a referral, Italian physicians must complete a standardized form, including a free-text field known as the Clinical Question, where they specify the reason for the prescription. The unstructured nature of this field, coupled with the absence of standardized alternatives, poses a major challenge for systematic analysis. Addressing this issue requires Natural Language Processing (NLP) techniques. In this study, we frame the task as a clustering problem: clinical questions are grouped using Transformer-based embeddings [6] to capture semantic similarity, and clusters are subsequently mapped to guideline-derived categories of appropriateness.
@@
Despite its importance, research on NLP-driven analysis of Italian referrals is limited. Venturelli et al. [7] used proprietary commercial software (Clinika VAP [8]) to classify referrals as appropriate or inappropriate, but without explicitly identifying the reasons for prescriptions. To the best of our knowledge, the only other study applying NLP on Italian referrals’ texts is [9], which focused on extracting and analyzing follow-up waiting times from clinical questions rather than assessing referrals appropriateness.
@@
Most previous studies on appropriateness in the Italian NHS relied on structured data or manual text analysis, focusing on medication prescriptions rather than medical examinations [10, 11, 12]. Research on the appropriateness of examinations is more limited [13, 14] and has not incorporated textual data. Other studies focused on the development of decision-support systems for guideline adherence, both in primary care [15, 16] and hospital settings [17, 18]. While effective in specific scenarios [19, 20], such tools are not suited for retrospective analyses of their performance or for quantifying the prevalence of appropriate or inappropriate referrals.
@@
Internationally, several studies have employed NLP and rule-based systems to assess referral appropriateness, often in disease-specific contexts. Haddad et al. [21] analyzed Russian referrals’ texts for appropriateness using labelled data; Sagheb et al. [22] designed a rule-based algorithm to verify guidelines adherence for asthma patients from EHR documents; similarly in [23] a rule-based approach was applied to assess guideline adherence for gout patients. Dimitrov et al. [24] conducted a nationwide evaluation of breast cancer guideline adherence in EHRs using proprietary commercial software (Danny platform [25]), which incorporated rules and manual reviews. Villeana et al. [26] analyzed Chilean referrals’ texts using labelled data to binary classify them for correct priority classes.
However, all these approaches rely on labeled data or proprietary software, limiting their adaptability and scalability across different examination types.
To address these gaps, this study proposes the first fully unsupervised NLP pipeline for analyzing appropriateness in Italian referrals. The pipeline is applicable to any examination for which guidelines define appropriate and inappropriate referral reasons. We developed this approach on a first case study related to venous echocolordoppler of the lower limbs (ECD) and we present results also on a second case study related to flexible endoscope colonoscopy (FEC). Both datasets include all referrals issued during three years in the entire the Lombardy Region, the most populated Italian region, with approximately 10 million inhabitants - akin to a small state.
These two examinations were selected because both are characterized by a high prevalence of inappropriate referrals [27, 28, 29, 30] and had already been targeted by Lombardy’s 2015 resolutions on angiology and gastroenterology [31, 32]. However, no comprehensive study has assessed the impact of those resolutions at scale. Our analysis provides, for the first time, a regional-scale evaluation of referral appropriateness and of guideline effectiveness, which informed the Lombardy Region’s December 2023 resolution [33].
This work thus provides both a methodological contribution, by introducing a scalable NLP pipeline for appropriateness assessment, and an applied contribution, by generating evidence that informed regional policies.
The remainder of this paper is structured as follows: Section 2 describes the datasets and the pipeline, Section 3 presents the results, Section 4 addresses implications, limitations, and future developments; and Section 5 concludes.
2 Methods
@@ 2.1 Data
2.1.1 Referrals datasets
The datasets analyzed in this study are extracted from all referrals issued in the Lombardy Region between 2019 and 2021 for ECD and FEC, totalling 496,971 referrals for ECD and 407,949 for FEC. Each referral includes the clinical question, the year, the Local Health Authority (LHA) of the prescribing physician and the type of prescribing physician. Physician types are categorized as general practitioners (GPs) or specialized physicians, while LHAs are the eight healthcare administrative areas within the Lombardy Region. Figure 1 illustrates the distribution of referrals among these groups.
Figure 1: Referrals distributions across years (A, B), type of prescribing physicians (C, D) and LHA (E, F) for ECD (A,C,E) and FEC (B,D,F).
The analysis focuses on the clinical question (CQ), a free-text field typically containing one or a few sentences. These texts exhibit frequent spelling and typing errors, heavy use of abbreviations, inconsistent punctuation, and non-standard capitalization. The dataset for ECD includes 229,509 distinct CQ, with only 15% appearing more than once, while the dataset for FEC includes 89,734 distinct CQ, with 41% appearing more than once. The median length for ECD is 71 characters (IQR=39), while for FEC it is 57 (IQR=20). Only 189 for ECD (0.08%) and 137 for FEC (0.15%) contain a meaningful ICD code, underscoring the necessity of employing NLP for this analysis.
The clinical questions do not contain personal information about patients or physicians.
@@ 2.1.2 Manual annotations
To validate the pipeline, a randomly selected subset of 1,000 distinct clinical questions was manually labelled for both ECD and FEC. The labels correspond to categories representing diseases or symptoms that prompted the referral. During an initial review of this subset, ambiguous cases were identified, and an angiologist and a gastroenterologist were consulted to resolve these doubts and align the clusters with prescription guidelines. Annotation guidelines were subsequently refined, and the entire subset was annotated accordingly. Two other reviews of the annotations corrected remaining errors and clarified uncertainties, ensuring the final labelled datasets were error-free. More details on the manual annotations are provided in Additional File 1, Section A.
@@ 2.1.3 Prescription guidelines
The guidelines used as a reference to assess referral appropriateness are extracted from [31] for ECD and from [32] for FEC. In particular, we extracted the appropriate and inappropriate diseases and symptoms, considering separately also cases that are more likely to be appropriate or more likely to be inappropriate, depending on details that might not be available in clinical questions. English schematized summaries of these guidelines are reported in Figure 2.
Figure 2: Guidelines for referrals for ECD and FEC in the Lombardy Region. Derived from the the Appendix 3 of [31] and the Appendix 2 of [32], translated into English.
2.2 Evaluation of prescription appropriateness
@@
Figure 3 depicts our pipeline with the blocks that compose it, and below, we detail each block.
@@ 2.2.1 Pre-processing
Physicians write referrals during patients’ examinations, with limited time available. This leads them to write clinical questions that are far from well-readable and correct texts. The grammar of the sentences is often poor, there are frequent typing errors, and the punctuation is either not used or used incorrectly. The use of uppercase letters does not have any particular meaning, and the use of abbreviations is very frequent. Even in the two types of examination analyzed in this study, multiple abbreviations can indicate the same disease/symptoms. In the manually labelled subsets, 15% of clinical questions were marked as having typing errors and 68% as containing abbreviations.
To address these issues, suitable pre-processing steps are applied:
@@
We remove all punctuation and lowercase sentences due to the inconsistent use of these elements.
For abbreviation expansion, we extract all the unique words with length $\leq~3$, and we identify those corresponding to abbreviations of medical terms, defining a vocabulary for their replacement in the texts.
To fix the typing mistakes, all words appearing with frequency $1\%$ and having another word similar to them with frequency $1\%$ are selected. These words are considered misspelt, and the most similar word with frequency $1\%$ is chosen as their correction. The similarity is measured with edit distance, and the minimum threshold is chosen by minimizing errors in a manually reviewed sample of 100 corrections.
Examples of pre-processing effects and an ablation analysis are reported in Additional File 1, Section B.
@@ 2.2.2 Text embedding
To appropriately cluster the clinical questions, it is necessary to produce a representation capturing the central semantics of these texts without getting confused by their widely present noise.
Considering the many synonyms and composite terms of the medical jargon used in these texts, vocabulary-based representations, such as TF-IDF [34], cannot be expected to work well for this task.
@@
On the contrary, context-based embeddings produced by transformers model consider the surrounding context while embedding each token and have often shown superior performances in the recent clinical NLP literature [36, 37, 38, 39]. The problem is particularly challenging in our case due to the limited number of embedding models available for the Italian language, especially for the clinical domain.
To this purpose, we adopt the same model developed in [40] to cluster sentences extracted from Italian discharge summaries, and we verify that even with referrals’ clinical questions, its performances are superior to other models. This model consists of a fine-tuned version of the last Italian version of BERT, Umberto [41], on a dataset of publicly available medical documents in Italian and other languages (translated into Italian) [42].
@@ 2.2.3 Clustering
Some studies have shown that HDBSCAN often outperforms other clustering algorithms—such as K-Means, hierarchical clustering, and DBSCAN-when applied to text embeddings, particularly when preceded by a dimensionality reduction step [43, 44]. Text embeddings typically lie on complex, non-linear manifolds and may form clusters with varying densities and irregular shapes. Algorithms like K-Means assume convex cluster geometries and require predefining the number of clusters, which is impractical for our task. Hierarchical clustering is sensitive to noise and outliers and lacks the ability to leave out unclustered data points, often producing forced or less meaningful groupings. In contrast, HDBSCAN does not require the number of clusters to be specified in advance and allows some data points to remain unclustered, making it a better fit for our scenario. Based on these advantages and on a comparison of the results on the ECD dataset, we selected HDBSCAN for the clustering step.
HDBSCAN requires specifying some hyperparameters, including the minimum cluster size, the minimum number of samples in the neighborhood of a core point (min_samplesmin\_samples), and the epsilon. To better accommodate the varying sizes of clusters in our dataset, we adopted a recursive strategy: the minimum cluster size was iteratively reduced, halving at each step, and applied only to data points that remained unclustered in the previous iteration. For the dimensionality reduction, we compared PCA, UMAP and no dimensionality reduction.
We performed a grid search and we selected the configuration with PCA with 80% of variance threshold, reducing the minimum cluster size from 500 to 10, with min_samples=5min\_samples=5 and $\epsilon=0.1$. The complete grid search results, together with comparisons to k-means and agglomerative hierarchical clustering, are reported in Additional File 1, Section C.
Due to resource constraints, it was not feasible to run HDBSCAN on the entire datasets. Therefore, we divided the data into chunks of 30,000 clinical questions and clustered them independently. In later stages, the summarization and mapping procedures allow to merge clusters from different chunks that correspond to the same diseases or symptoms.
@@ 2.2.4 Cluster summarization
Summarizing the identified clusters is necessary to understand the symptoms or diseases corresponding to each cluster. To automatize this step, we extracted keywords from each cluster. The most frequent words in each cluster were initially selected, considering a decreasing frequency threshold from 80% until at least two keywords were found. Among these, we excluded those with high frequency in at least half of the clusters. This allowed us to exclude not only stop words but also frequent words for the specific examination we are analyzing.